# Obstacle-Avoidant Leader Following with a Quadruped Robot

### v1

@@ Abstract
Personal mobile robotic assistants are expected to find wide applications in industry and healthcare. For example, people with limited mobility can benefit from robots helping with daily tasks, or construction workers can have robots perform precision monitoring tasks on-site. However, manually steering a robot while in motion requires significant concentration from the operator, especially in tight or crowded spaces. This reduces walking speed, and the constant need for vigilance increases fatigue and, thus, the risk of accidents. This work presents a virtual leash with which a robot can naturally follow an operator. We use a sensor fusion based on a custom-built RF transponder, RGB cameras, and a LiDAR. In addition, we customize a local avoidance planner for legged platforms, which enables us to navigate dynamic and narrow environments. We successfully validate on the ANYmal platform [1] the robustness and performance of our entire pipeline in real-world experiments.
@@ I Introduction
Most existing human-following solutions for mobile ground robots either rely on tethered connections between the robot and user (i.e., a physical leash) [6], use sensors to track a transmitter on the person [7, 8, 9], or follow a human by using a range measurement sensor and maintaining a constant distance to the nearest obstacle in front of them, predicated on the assumption that this is the human [10, 11].
@@ II Related Work
Few previous works exist on ground robots that actively track and follow a human could be found. Peña et al. [13] developed a human-following robot using a camera to track a triangle formation of three IR transmitters on the target’s back instead of the human directly. This approach is intrusive and perspective-limited (tracking is only possible from behind). Li et al. [9] use a quadrupedal robot to track a beacon, using externally mounted base stations on-site. Their approach limits the deployment range to the area spanned by the beacons, severely limiting deploy- and scalability. Having all required sensors mounted directly onto the robot lets us directly overcome such limitations. Several image-based leader tracking approaches exist for aerial vehicles where issues on perspective, line of sight, and path planning are much simplified [2, 3]. Commercial approaches to leader tracking on a legged robot [14] and aerial vehicles [4] exist, showcasing their relevance. However, no information exists on the underlying implementation.
@@ III-B Leader tracking
Inspired by literature, we track our leader using an Extended Kalman Filter (EKF), which uses a linear motion model. To initialize the EKF, we wait for two consistent leader hypotheses to filter out outliers. From these two initial matches, we derive an initial guess on the leader velocity in the $xy$ plane of the robot. The resulting leader parametrization, consisting of a position $p_{0}^{l}\in\mathbb{R}^{2}$, velocity $v_{0}^{l}\in\mathbb{R}^{2}$, and a set of 3D LiDAR points on the leader $P_{0}^{l}\in\mathbb{R}^{n\times 3}$, are used to initialize leader tracking. We use the prediction step from the EKF together with the motion estimation of the robot to project the previously detected leader points $P^{l}_{t-1}$ into the current time frame to obtain the predicted leader point cloud $\hat{P}^{l}_{t}$. Before matching them with the latest LiDAR scan $P_{t}$, we remove large plains (i.e., walls and the floor) using RANSAC [27] to eliminate 3D structures we are certain are not humans. We then select new leader points through a KD-tree nearest-neighbor search [28] by matching each point in $\hat{P}^{l}_{t}$ with the nearest points in $P_{t}$. After each search association step, we reject 40% of the points furthest from the predicted leader center to obtain the final measured $P^{l}_{t}$. This heuristic prevents the leader point cloud from spreading onto adjacent static objects or people while not degrading the performance of the EKF. When available, we fuse the image segmentation to bound the association step. When the leader is occluded or out of view of the cameras, we rely solely on the EKF motion model.
@@ IV-A Hardware
Figure 4: We modify waverider to track our leader while autonomously avoiding static obstacles, such as benches and door frames. We create a wavemap occupancy map, actively filtering out human points to only include the static scene. The map resulting from a tracking sequence can be seen in (A). Said sequence is visualized in (B), with only the final leader frame visualized to reduce visual clutter. In (C), we show the path the robot takes while avoiding collisions. The leader trajectory (as recorded by the robot) is visualized via the purple line and the robot trajectory via the sequence of arrow clusters. Each cluster represents a pose and the forces acting on the robot during that single time step, the green symbolizing the attractive force of the goal and the red the repelling forces of the obstacles within range. The image is split into three frames to highlight the multi-resolution nature of waverider, with the green blocks representing the obstacles. Each third of the image represents one timestep, and the black cross shows the exact robot position for that timestep.

### v2

@@ Abstract
Personal mobile robotic assistants are expected to find wide applications in industry and healthcare. For example, people with limited mobility can benefit from robots helping with daily tasks, or construction workers can have robots perform precision monitoring tasks on-site. However, manually steering a robot while in motion requires significant concentration from the operator, especially in tight or crowded spaces. This reduces walking speed, and the constant need for vigilance increases fatigue and, thus, the risk of accidents. This work presents a virtual leash with which a robot can naturally follow an operator. We use a sensor fusion based on a custom-built RF transponder, RGB cameras, and a LiDAR. In addition, we customize a local avoidance planner for legged platforms, which enables us to navigate dynamic and narrow environments. We successfully validate on the ANYmal platform [1] the robustness and performance of our entire pipeline in real-world experiments. The video is available at: obstacle-avoidant-leader-following.
@@ I Introduction
Most existing human-following solutions for mobile ground robots either rely on tethered connections between the robot and user (i.e., a physical leash) [6], use sensors to track a transmitter on the person [7, 8, 9], use purely image-based tracking [10], or follow a human by using a range measurement sensor and maintaining a constant distance to the nearest obstacle in front of them, predicated on the assumption that this is their human [11, 12].
@@ II Related Work
Few previous works exist on ground robots that actively track and follow a human could be found. Peña et al. [14] developed a human-following robot using a camera to track a triangle formation of three IR transmitters on the target’s back instead of the human directly. This approach is intrusive and perspective-limited (tracking is only possible from behind). Li et al. [9] use a quadrupedal robot to track a beacon, using externally mounted base stations on-site. Their approach limits the deployment range to the area spanned by the beacons, severely limiting deployability and scalability. Having all required sensors mounted directly onto the robot allows us to overcome such limitations directly. A purely image-based leader tracking was proposed by Rollo et al. [10], however, it was not shown to work in dynamic or crowded environments. Several image-based leader tracking approaches exist for aerial vehicles where issues on perspective, line of sight, and path planning are much simplified [2, 3]. Commercial approaches to leader tracking on a legged robot [15] and aerial vehicles [4] exist, showcasing their relevance. However, no information exists on the underlying implementation.
@@ III-B Leader Tracking
Inspired by literature, we track our leader using an Extended Kalman Filter (EKF), which uses a linear motion model. To initialize the EKF, we wait for two consistent leader hypotheses to filter out outliers. From these two initial matches, we derive an initial guess on the leader velocity in the $xy$ plane of the robot. The resulting leader parametrization, consisting of a position $p_{0}^{l}\in\mathbb{R}^{2}$, velocity $v_{0}^{l}\in\mathbb{R}^{2}$, and a set of 3D LiDAR points on the leader $P_{0}^{l}\in\mathbb{R}^{n\times 3}$, are used to initialize leader tracking. We use the prediction step from the EKF together with the motion estimation of the robot to project the previously detected leader points $P^{l}_{t-1}$ into the current time frame to obtain the predicted leader point cloud $\hat{P}^{l}_{t}$. Before matching them with the latest LiDAR scan $P_{t}$, we remove large planes (i.e., walls and the floor) using RANSAC [28] to eliminate 3D structures we are certain are not humans. We then select new leader points through a KD-tree nearest-neighbor search [29] by matching each point in $\hat{P}^{l}_{t}$ with the nearest points in $P_{t}$. After each search association step, we reject 40% of the points furthest from the predicted leader center to obtain the final measured $P^{l}_{t}$. This heuristic prevents the leader point cloud from spreading onto adjacent static objects or people while not degrading the performance of the EKF. When available, we fuse the image segmentation to bound the association step. When the leader is occluded or out of view of the cameras, we rely solely on the EKF motion model.
@@ IV-A Hardware
Figure 4: We modify waverider to track our leader while autonomously avoiding static obstacles, such as benches and door frames. We create an occupancy map with wavemap, actively filtering out human points to only include the static scene. The map resulting from a tracking sequence can be seen in (A). Said sequence is visualized in (B), with only the final leader frame visualized to reduce visual clutter. In (C), we show the path the robot takes while avoiding collisions. The leader trajectory (as recorded by the robot) is visualized via the purple line and the robot trajectory via the sequence of arrow clusters. Each cluster represents a pose and the forces acting on the robot during that single time step, the green symbolizing the attractive force of the goal and the red the repelling forces of the obstacles within range. The image is split into three frames to highlight the multi-resolution nature of waverider, with the green blocks representing the obstacles. Each third of the image represents one timestep, and the black cross shows the exact robot position for that timestep.